The length of the base of a rectangular prism is twice its width. The perimeter of the base is 42 yards. If the volume of the prism is 980 cubic yards, find the height of the prism.


1
Expert's answer
2022-03-21T16:06:11-0400

Width = x

Length = 2x

Perimeter = x + 2x + x + 2x = 42

6x = 42; x = 42/6

x = 7

Width = 7 yards

Length = 2*7 = 14 yards

Volume = whl = 980

7 * h * 14 = 980

98h = 980; h = 980/98

h = 10 yards